OpenStudy (anonymous):

(x – 16)2 + (y + 9)2 = 4 ?

OpenStudy (anonymous):

y = 9-x and i got that by FOILing

OpenStudy (chaise):

What do you want to know? It's a circle with origin 16,9 and radius 2. DoomDude, I think this is in actual fact: (x – 16)^2 + (y + 9)^2 = 4 ?

OpenStudy (anonymous):

(x – 5)2 + (y – 1)2 = 25?

OpenStudy (chaise):

Radius: 5 Origin: (5,1)

OpenStudy (anonymous):

Circle A has an equation of (x + 8)2 + (y + 3)2 = 16. Where is the center of the circle located?

OpenStudy (chaise):

(-8,-3)

OpenStudy (anonymous):

Circle T has an equation of x2 + (y – 10)2 = 25?

OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, 1). Which equation defines circle P?

OpenStudy (anonymous):

can someone please help me with Circle P has a radius of 8 units with center P at (2, 1). Which equation defines circle P?